The region above the x-axis and below the curve 𝑦 = 𝑒 βˆ’π‘₯ π‘“π‘œπ‘Ÿ 0 ≀ π‘₯ ≀ ∞ is rotated around the x-axis. Find the volume of the solid generated.


1
Expert's answer
2021-06-10T09:56:09-0400

"V=|\int^{\infin}_0\pi(e^{-x})^2dx|=|\frac{\pi}{2e^{2x}}|^{\infin}_0|=\pi/2"